OpenStudy (anonymous):

Combine and simplify the following radical expressions. \[6\sqrt[3]{-16}+\frac{ 2 }{ 3 } \sqrt[3]{-54}-5\sqrt[3]{250}\]

OpenStudy (anonymous):

I'm doing a review and its been a long time since I've done these. If anyone can guide me through the steps that would be great.

OpenStudy (perl):

I would start by factoring the expressions under the radical sign

OpenStudy (jdoe0001):

\(\large { 6\sqrt[3]{-16}+\frac{ 2 }{ 3 } \sqrt[3]{-54}-5\sqrt[3]{250}\qquad \begin{cases} 16\to 2\cdot 2\cdot 2\cdot 2\to &2\cdot 2^3\\ 54\to 2\cdot 3\cdot 3\cdot 3\to &2\cdot 3^3\\ 250\to 2\cdot 5\cdot 5\cdot 5\to& 2\cdot 5^3 \end{cases} \\ \quad \\ 6\sqrt[3]{-16}+\cfrac{ 2 }{ 3 } \sqrt[3]{-54}-5\sqrt[3]{250}\implies ? }\)

OpenStudy (anonymous):

@perl Is this correct for after I factor it? \[12\sqrt[3]{2}+2\sqrt[3]{-2}-25\sqrt[3]{-2}\]

OpenStudy (anonymous):

Whoops put negatives in wrong. \[12\sqrt[3]{-2}+2\sqrt[3]{-2}-25\sqrt[3]{2}\]

OpenStudy (jdoe0001):

\(\large { 6\sqrt[3]{-16}+\frac{ 2 }{ 3 } \sqrt[3]{-54}-5\sqrt[3]{250}\qquad \begin{cases} 16\to 2\cdot 2\cdot 2\cdot 2\to &2\cdot 2^3\\ 54\to 2\cdot 3\cdot 3\cdot 3\to &2\cdot 3^3\\ 250\to 2\cdot 5\cdot 5\cdot 5\to& 2\cdot 5^3 \end{cases} \\ \quad \\ 6\sqrt[3]{-16}+\cfrac{ 2 }{ 3 } \sqrt[3]{-54}-5\sqrt[3]{250} \\ \quad \\ 6\sqrt[3]{-2^3\cdot 2}+\cfrac{ 2 }{ 3 }\sqrt[3]{-3^3\cdot 2}-5\sqrt[3]{-5^3\cdot 2} \\ \quad \\ \begin{cases} -2\cdot -2\cdot -2\to -2^3\\ -3\cdot -3\cdot -3\to -3^3\\ -5\cdot -5\cdot -5\to -5^3 \end{cases} \\ \quad \\ 6\sqrt[3]{-2^3\cdot 2}+\cfrac{ 2 }{ 3 }\sqrt[3]{-3^3\cdot 2}-5\sqrt[3]{-5^3\cdot 2} \\ \quad \\ 6\sqrt[3]{{\color{brown}{ -2}}^3} \cdot \sqrt[3]{2}+\cfrac{ 2 }{ 3 }\sqrt[3]{{\color{brown}{ -3}}^3}\sqrt[3]{2}-5\sqrt[3]{{\color{brown}{ -5}}^3} \cdot \sqrt[3]{ 2} \\ \quad \\ 6\cdot {\color{brown}{ -2}}\sqrt[3]{ 2}+\cfrac{ 2 }{ 3 }\cdot {\color{brown}{ -3}}\sqrt[3]{ 2}-5\cdot {\color{brown}{ -5}}\sqrt[3]{ 2}\implies ? }\)